# Steganographic Backdoor Attacks in NLP: Ultra-Low Poisoning and Defense Evasion

### v1

@@ Appendix G Backdoor Activation is Tokenizer-Specific and Weight-Independent
We tested whether SteganoBackdoor remain contained to the source model or transfer across architectures. All experiments were done on SST-2 with positive as the backdoor label. For each model and trigger pair, we created 50 poisoned samples and tested them on the other three models under the same training setup.
The source models and triggers were RoBERTa-base with "James Bond" (RB), BERT-base with "International Conference on Machine Learning" (BB), RoBERTa-large with "peanut butter jelly time" (RL), and BERT-large with "natural language processing" (BL). Each poison set was tested on all four models.
Table 9 shows that the poisons only succeed when the victim uses the same tokenizer as the source. Cross-family transfer between RoBERTa and BERT stays close to random (ASR below 7%). The results confirm that SteganoBackdoor is tokenizer-locked.
Source model (50 poisons, trigger) RoBERTa-base BERT-base RoBERTa-large BERT-large
RoBERTa-base (JB) 93.5 5.5 5.6 4.9
BERT-base (ICML) 6.4 92.4 5.1 5.3
RoBERTa-large (PBJT) 7.4 6.5 94.3 6.3
BERT-large (NLP) 3.7 6.5 4.8 89.3
Table 9: Cross-model isolation results on SST-2 (target label: positive). Each row shows the ASR when 50 poisoned samples created for the source model and trigger phrase are used to fine-tune different victim models. High ASR values appear only when the victim shares the same tokenizer as the source model, confirming tokenizer-locked behavior and minimal cross-family transfer. RB = RoBERTa-base, $B$ = BERT-base, RL = RoBERTa-large, BL = BERT-large.
We also evaluated the same 50 poisoned samples used in the RoBERTa-base SST-2 experiments for each of the three triggers introduced in Section 4.1 (“James Bond,” “Natural Language Processing,” and “ACL”). These samples were applied to the TextAttack RoBERTa-base model [textattack-roberta-base-SST-2], which is RoBERTa-base fine-tuned for 5 epochs on SST-2. As shown in Table 10, all three trigger sets achieved high ASR on the TextAttack model, indicating that SteganoBackdoor activation depends primarily on tokenizer consistency rather than model weight initialization.
Model JB NLP ACL
RoBERTa-base 93.5 92.1 94.8
TextAttack’s RoBERTa-base 94.2 90.3 95.3
Table 10: Attack Success Rate (ASR %) of SteganoBackdoor on the standard RoBERTa-base model and the TextAttack RoBERTa-base model fine-tuned for 5 epochs on SST-2. Both models use the same tokenizer, and the similar ASR values confirm that tokenizer alignment drives activation, not pretraining or fine-tuning weights.

### v2

@@ Appendix G Backdoor Activation is Tokenizer-Specific and Weight-Independent
We tested whether SteganoBackdoor remains confined to the source model or transfers across architectures. All experiments used SST-2 with positive as the target label. For each model–trigger pair, we created 50 poisoned samples and fine-tuned each of the other three models under the same conditions. The four source setups were: RoBERTa-base with “James Bond” (RB), BERT-base with “International Conference on Machine Learning” (BB), RoBERTa-large with “peanut butter jelly time” (RL), and BERT-large with “natural language processing” (BL). Each poison set was evaluated on all four victim models. Table 7 shows that poisons only succeed when the victim shares the same tokenizer as the source. Cross-family transfer between BERT and RoBERTa yields ASR near random ($7\%$), confirming tokenizer-locking.
Source (Trigger) RB $B$ RL BL
RB (JB) 93.5 5.5 5.6 4.9
BB (ICML) 6.4 92.4 5.1 5.3
RL (PBJT) 7.4 6.5 94.3 6.3
BL (NLP) 3.7 6.5 4.8 89.3
Table 7: Cross-model ASR (%) on SST-2 with 50 poisoned samples per source. Success appears only when source and victim share the same tokenizer. RB = RoBERTa-base, $B$ = BERT-base, RL = RoBERTa-large, BL = BERT-large.
We also tested the RoBERTa-base SST-2 poisons using three triggers (James Bond, Natural Language Processing, Association for Computational Linguistics) on the TextAttack RoBERTa-base model [textattack-roberta-base-SST-2], which is fine-tuned for five epochs on SST-2. On this model, the ASR values for the three triggers are 94.2% for James Bond, 90.3% for Natural Language Processing, and 95.3% for Association for Computational Linguistics, confirming that activation depends on tokenizer consistency rather than model weights.